OpenStudy (anonymous):

Find \[\log_{x ^{5}} y \times \log_{y} x ^{2}\] where x, y > 0

OpenStudy (cwtan):

logx5 y=1/logy x^5 logy x^5=5 logy x and also logy x^2=2logy x so when they multiply togther 2logy x/5 logy x=2/5

OpenStudy (anonymous):

Thank you